solve the eqation y+4/5=7/8
Found 2 solutions by richwmiller, checkley79:
Answer by richwmiller(17219)

convert 4/5 to x/40 4/5=32/40
convert 7/8 to y/40 7/8=35/40
y+32/40=35/40
y=3/40

Answer by checkley79(3341)

y+4/5=7/8 subtract 4/5 from both sides.
y=7/8-4/5 convert fractions to common denominators (40).
y=(7*5)/40-(4*8)/40
y=35/40-32/40 subtract numerators.
y=3/40 answer.